OpenStudy (anonymous):

Which is the area of a triangle with a base of 10 inches and a height of 3 inches?Select one of the options below as your answer:A. 15 square inchesB. 24 square inches C. 48 square inches D. none of the above

OpenStudy (hoblos):

3*10/2 = 30/2 =15

OpenStudy (anonymous):

ty

OpenStudy (hoblos):

yw